OpenStudy (anonymous):

Given f(x) and g(x) = f(x) + k, look at the graph below and determine the value of k. k = _______

OpenStudy (anonymous):

Graph:

OpenStudy (anonymous):

k=4

OpenStudy (anonymous):

I'm not entirely sure I trust you... what with the prior harassment and everything... >_>

OpenStudy (anonymous):

I wouldn't put it past the epic 10/10 trollololo mastur of 2013 to give me an incorrect answer. lol

OpenStudy (anonymous):

LOL well i can explain if you'd like

OpenStudy (anonymous):

Okay, yeah, sure.

OpenStudy (anonymous):

basically its saying this equation g(x) is equal to f(x) + some number. If you look at the graph they are two different lines. To make them g(x) and f(x) the same line what number would you move f(x) up by ? And well if you count its 4 units. Therefore k=4. Or you can do it the math way which I'll post in a sec

OpenStudy (anonymous):

so g(x) = -1/3x+1 f(x) =-1/3x -3 if g(x) = f(x)+ k -1/3x+1 = -1/3x-3+k 1=-3+k k=4

OpenStudy (anonymous):

swaggggggooooo.

OpenStudy (anonymous):

Ohh, okay. I get it now. It was simpler than I thought... I guess I was over-thinking it. Thanks for your help; I appreciate it.

OpenStudy (anonymous):

no problem. lI'll go back to 4chan whatever that is

OpenStudy (anonymous):

Lol, okay~ o3o

OpenStudy (anonymous):

jks i dont even know what 4chan is. But I do have to get back to discrete math. q.q wanna help me with that? XD